how to solve a problem in which there are two real numbers and taking half of one number is the same as adding 12.5 to another. Also the squares of the two real numbers are 200 apart?
Answer by stanbon(75887)

how to solve a problem in which there are two real numbers and taking half of one number is the same as adding 12.5 to another. Also the squares of the two real numbers are 200 apart?
-------
Equations::
(1/2)x = y+12.5
--------------------
x^2 - y^2 = 200
----------------------
x = 2y+25
Substitute for "x" and solve for "y"::
(2y+25)^2 - y^2 = 200
--------
3y^2 + 100y + 425 = 0
----
y = -5
----
x^2 = 200-(5)^2 = 175
x = 13.23
------------------------------
Cheers,
Stan H.
-------------------